# Transforming Location Retrieval at Airbnb: A Journey from Heuristics to Reinforcement Learning

### v1

@@ Abstract.
copyright: noneccs: Information systems Recommender systemsccs: Information systems Novelty in information retrievalccs: Information systems Query intentccs: Applied computing Online shopping
@@ 1. Introduction
 Airbnb has incredible scale with over 7M+ active listings, $\sim$390M nights booked per year and proportionately large search volume. It also has complex ranking systems with strict performance and product expectations. As a result, it is infeasible to rank every listing for every search. We also cannot simply filter by whether a listing matches the exact searched location, due to the value and diverse amenities of listings in surrounding areas discussed above. Our search system requires defining a rectangular area on the map, known as retrieval bounds. These retrieval bounds are used to filter our inventory to a reasonable set of listings that any guest would ever book for the search. The search ranking models described in (Abdool et al., 2020) (Haldar et al., 2019) (Haldar et al., 2020) (Tan et al., 2023) (Haldar et al., 2023) then determine the most relevant and bookable listings in these retrieval bounds, which are shown to Airbnb’s guests. This flow is outlined below in Figure 2
Figure 2. Location Retrieval System Architecture
@@
There is little existing work on how to tackle the location retrieval problem even though location based search is used across the web every day on products like Yelp™, Google Maps™, Booking.com™, and Zillow™. The closest task is audience expansion for applications like marketing and ads which have been explored by many (Liu et al., 2016) (Zhuang et al., 2020) (Chan et al., 2021). However, none of their approaches could directly be applied due to the large differences in task construction and output such as optimizing similarity instead of diversity and product experience constraints regarding the distance of listings. This paper will go over the evolution and efficacy of solutions for the outlined challenges, including simple heuristics, statistics, machine learning and reinforcement learning, with an emphasis on the application of reinforcement learning.
@@ 4.1. The opportunity
 For example, large group travelers searching for San Francisco book in significantly different locations than San Francisco searchers overall, demonstrated in Figure 1 and 3. Differentiating search experiences based on parameters like guest count and lead time could potentially provide more succinct, relevant results for guests.
@@ 5.4. Uncertainty Estimation
    Ridge regression (Hoerl and Kennard, 1970): regularized linear regression with least squares loss that allows for uncertainty estimation based on the standard errors of its coefficients
@@
    Bayesian deep learning (Neal, 2012)(Joo et al., 2020): Deep neural networks where weights and biases are constructed as random variables with distributions for uncertainty estimation.
@@
    Deep ensembles (Lakshminarayanan et al., 2017): Ensemble of deep neural networks that are jointly use to compute the prediction mean and variance that can be used for uncertainty estimation.
@@
    Evidential Networks (Sensoy et al., 2018): Utilizes learned belief functions to model the predictors uncertainty
@@
 Given these considerations, we employ Monte Carlo Dropout(Gal and Ghahramani, 2016) (Guo et al., 2020) to handle uncertainty estimation. Gal et. al demonstrates that any deep neural network of arbitrary depth, non linearities, and architecture with a dropout layer is equivalent to an approximation to the probabilistic deep Gaussian process. Given an example, we can score it N unique times to generate N unique predictions which is equivalent to the score distribution for the example. This score distribution can be used for uncertainty estimation. This approach requires minimal modification to the existing model architecture and any additional overhead for training and serving can be minimized through parallelization unlike some of the other methods.
@@ 5.5. Exploration Strategy
    E-greedy(Sutton and Barto, 2018): Greedily recommends the actions with the highest reward with probability 1-e and randomly selects other actions uniformly with probability e.
@@
    Thompson sampling(Agrawal and Goyal, 2013): Samples from the posterior distribution of the action reward space with current knowledge; recommends the action with the highest sampled reward estimate.
@@
    Upper Confidence Bound(Auer, 2002): Computes upper confidence bounds for each action using the mean estimate for an action plus a bonus based on the uncertainty of the estimate and tuneable parameter $\lambda$. It then chooses the action with the highest upper confidence bound.
@@ 6.1. Overview
Reinforcement Learning Machine Learning +0.51% +0.25% -9.3% +0.01%
@@ 6.2. Cold Start Heuristic based Location Retrieval
 The log scale expansion function (Heuristic 4) was tested in online A/B experiments against the baseline method (Heuristic 3) outlined in Section 2 for each location type. It significantly improved the launch criterion of uncancelled bookers by 0.35% and was launched to 100% of production traffic. Offline impact is not shown in Table 1 because it was developed and tested before our offline evaluation criteria was defined and tracked.
 These improvements proved that location retrieval was an impactful lever to improve experiences for guests and booking behavior. These heuristics improved searches that had disproportionately low inventory states and unbookable inventory. For example searches for a specific building in Ibiza, Spain previously resulted in 5 results with the baseline while the new heuristic surfaced several pages of bookable results by expanding the bounds less than a mile. This is demonstrated in Figures 5 and 6.
Figure 5. Search for Building in Ibiza, Spain Before Log Scale Expansion Heuristic
Figure 6. Search for Building in Ibiza, Spain After Log Scale Expansion Heuristic
@@ 6.3. Statistics based Location Retrieval
 The statistics based method was tested in an online A/B experiment against the three heuristics for each searched location typed defined in Section 1. However, it did not have a stat sig impact on uncancelled bookers so it was never launched to 100% of production traffic. This experiment showed that naively taking advantage of booking data segmented only by searched location was not enough to materially improve the amount of bookers on the platform. We believed this was likely due to the methods inability to generalize across different searches and differentiate guest experiences because it was coarsely keyed by searched location. The need for this is evidenced by the difference in booking behavior between different guest counts, demonstrated in Figure 1 and 3 above.
@@ 6.4. Machine Learning based Location Retrieval
 The machine learning based location retrieval solutions were also tested successively in online A/B experiments resulting in a cumulative +1.8% uncancelled bookers by showing guests more bookable inventory near their travel destinations. Each iteration had statistically significant positive impact to uncancelled bookers, was launched to 100% of production traffic for a period of time, and served as the following iteration’s baseline for offline and online evaluation.
 We also qualitatively evaluated searches to verify whether the model generalizes for uncommon locations. For example, the baseline retrieval bounds are arbitrarily large with many results that were extremely far away and irrelevant for a search for a small street in Lima, Peru such that the searched street is not visible in the UI. The machine learning based retrieval bounds results in much smaller bounds and more relevant results near the street demonstrating its ability to generalize for uncommon locations. This is shown in Figures 7 and 8.
Figure 7. Retrieval Bounds and results for a search for a small street in Lima, Peru before machine learning based location retrieval
Figure 8. Retrieval Bounds and results for a search for a small street in Lima, Peru after machine learning based location retrieval
@@ 6.5. Reinforcement Learning based Location Retrieval
 The MC Dropout Location Retrieval model performs better than the baseline machine learning model when evaluating offline using the mean prediction $\mu$ for each test example with +0.25% in recall and -9.3% in retrieval size. We also qualitatively verified whether the MC Dropout Location Retrieval model explores more for locations with little prior data versus locations that are often searched and booked. For example, the difference between the mean bounds and UCB bounds for San Francisco, California is very small. However, the difference is much larger for Smith Mountain Lake, Virginia where we have little prior data. This is demonstrated in Figure 9.
Figure 9. Mean and UCB Retrieval Bounds from Searches San Francicso, CA (left) and Smith Mountain Lake, Virginia (right)
 The MC Dropout Location Retrieval Model was also tested in an online A/B experiment against the baseline production machine learning model retrained on data up to the start of the experiment and the production model trained on data up to 8 months before the start of the experiment. There was no statistically significant difference in performance between the production and retrained production models. The MC Dropout model showed a statistically significant increase of +0.51% in uncancelled bookers compared to the retrained production model, which was primarily driven by surfacing more bookable, affordable, and diverse inventory to guests. This is the result of two successive experiments. The first tested the method described against the preexisting machine learning solution. The second optimized the 32 forward passes required for uncertainty estimation by embedding it into the tensorflow graph instead of manually scoring the model 32 separate times which resulted in a significant latency improvement. Traditional multi-armed bandit regret analysis is infeasible for the nontraditional recall maximizing application but model confidence for each prediction offset has increased by 7.33% on average since launch.
@@ 7. Conclusion
We presented how location retrieval was built from the ground up in Airbnb search by solving cold start, generalization, and algorithmic bias. The system evolved from heuristics all the way to the application of reinforcement learning with significant cumulative impact of +2.66% in uncancelled bookers. This is comparable or more impactful than many previously published machine learning techniques within Airbnb search such as (Abdool et al., 2020) (Haldar et al., 2019) (Haldar et al., 2020) (Tan et al., 2023) (Haldar et al., 2023). We outlined the successes and failures along the way and the motivation for more advanced techniques to tackle the aforementioned challenges. The Monte Carlo Dropout model, the application of reinforcement learning, proved that it performed better than previous baselines offline and online for business metrics and now serves all production traffic at Airbnb. For future work, there is opportunity to advance the model’s understanding with more complex features. There is also an opportunity to reformulate the retrieval mechanism from retrieval bounds to map cells. This would allow for organically learning booking probabilities for each cell instead of training with hand tuned weights.

### v2

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USAbooktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAdoi: 10.1145/3627673.3680089isbn: 979-8-4007-0436-9/24/10ccs: Information systems Recommender systemsccs: Information systems Novelty in information retrievalccs: Information systems Query intentccs: Applied computing Online shopping
@@ 1. Introduction
 Airbnb has incredible scale with over 7M+ active listings, $\sim$390M nights booked per year and proportionately large search volume. It also has complex ranking systems with strict performance and product expectations. As a result, it is infeasible to rank every listing for every search. We also cannot simply filter by whether a listing matches the exact searched location, due to the value and diverse amenities of listings in surrounding areas discussed above. Our search system requires defining a rectangular area on the map, known as retrieval bounds. These retrieval bounds are used to filter our inventory to a reasonable set of listings that any guest would ever book for the search. The search ranking models described in (abdool2020managing) (haldar2019applying) (haldar2020improving) (tan2023optimizing) (haldar2023learning) then determine the most relevant and bookable listings in these retrieval bounds, which are shown to Airbnb’s guests.
@@
There is little existing work on how to tackle the location retrieval problem even though location based search is used across the web every day on products like Yelp™, Google Maps™, Booking.com™, and Zillow™. The closest task is audience expansion for applications like marketing and ads which have been explored by many (10.1145/2939672.2939680) (10.1145/3394486.3403295) (10.1145/3447548.3467179). However, none of their approaches could directly be applied due to the large differences in task construction and output such as optimizing similarity instead of diversity and product experience constraints regarding the distance of listings. This paper will go over the evolution and efficacy of solutions for the outlined challenges, including simple heuristics, statistics, machine learning and reinforcement learning, with an emphasis on the application of reinforcement learning.
@@ 4.1. The opportunity
 For example, large group travelers searching for San Francisco book in significantly different locations than San Francisco searchers overall, demonstrated in Figure 1 and 2. Differentiating search experiences based on parameters like guest count and lead time could potentially provide more succinct, relevant results for guests.
@@ 5.4. Uncertainty Estimation
    Ridge regression (hoerl1970ridge): regularized linear regression with least squares loss that allows for uncertainty estimation based on the standard errors of its coefficients
@@
    Bayesian deep learning (neal2012bayesian)(joo2020being): Deep neural networks where weights and biases are constructed as random variables with distributions for uncertainty estimation.
@@
    Deep ensembles (lakshminarayanan2017simple): Ensemble of deep neural networks that are jointly use to compute the prediction mean and variance that can be used for uncertainty estimation.
@@
    Evidential Networks (sensoy2018evidential): Utilizes learned belief functions to model the predictors uncertainty
@@
 Given these considerations, we employ Monte Carlo Dropout(gal2016dropout) (guo2020deep) to handle uncertainty estimation. Gal et. al demonstrates that any deep neural network of arbitrary depth, non linearities, and architecture with a dropout layer is equivalent to an approximation to the probabilistic deep Gaussian process. Given an example, we can score it N unique times to generate N unique predictions which is equivalent to the score distribution for the example. This score distribution can be used for uncertainty estimation. This approach requires minimal modification to the existing model architecture and any additional overhead for training and serving can be minimized through parallelization unlike some of the other methods.
@@ 5.5. Exploration Strategy
    E-greedy(sutton2018reinforcement): Greedily recommends the actions with the highest reward with probability 1-e and randomly selects other actions uniformly with probability e.
@@
    Thompson sampling(agrawal2013thompson): Samples from the posterior distribution of the action reward space with current knowledge; recommends the action with the highest sampled reward estimate.
@@
    Upper Confidence Bound(auer2002using): Computes upper confidence bounds for each action using the mean estimate for an action plus a bonus based on the uncertainty of the estimate and tuneable parameter $\lambda$. It then chooses the action with the highest upper confidence bound.
@@ 6.1. Overview
Reinforcement Learning Machine Learning +0.26% +0.25% -9.3% +0.01%
@@ 6.2. Cold Start Heuristic based Location Retrieval
 The log scale expansion function (Heuristic 4) was tested in online A/B experiments against the baseline method (Heuristic 3) outlined in Section 2 for each location type. It significantly improved the launch criterion of uncancelled bookers by 0.35% and was launched to 100% of production traffic. Offline impact is not shown in Table 1 because it was developed and tested before our offline evaluation criteria was defined and tracked.  These improvements proved that location retrieval was an impactful lever to improve experiences for guests and booking behavior. These heuristics improved searches that had disproportionately low inventory states and unbookable inventory. For example searches for a specific building in Ibiza, Spain previously resulted in 5 results with the baseline while the new heuristic surfaced several pages of bookable results by expanding the bounds less than a mile.
@@ 6.3. Statistics based Location Retrieval
 The statistics based method was tested in an online A/B experiment against the three heuristics for each searched location typed defined in Section 1. However, it did not have a stat sig impact on uncancelled bookers so it was never launched to 100% of production traffic. This experiment showed that naively taking advantage of booking data segmented only by searched location was not enough to materially improve the amount of bookers on the platform. We believed this was likely due to the methods inability to generalize across different searches and differentiate guest experiences because it was coarsely keyed by searched location. The need for this is evidenced by the difference in booking behavior between different guest counts, demonstrated in Figure 1 and 2 above.
@@ 6.4. Machine Learning based Location Retrieval
 The machine learning based location retrieval solutions were also tested successively in online A/B experiments resulting in a cumulative +1.8% uncancelled bookers by showing guests more bookable inventory near their travel destinations. Each iteration had statistically significant positive impact to uncancelled bookers, was launched to 100% of production traffic for a period of time, and served as the following iteration’s baseline for offline and online evaluation. We also qualitatively evaluated searches to verify whether the model generalizes for uncommon locations. For example, the baseline retrieval bounds are arbitrarily large with many results that were extremely far away and irrelevant for a search for a small street in Lima, Peru such that the searched street is not visible in the UI. The machine learning based retrieval bounds results in much smaller bounds and more relevant results near the street. This demonstrates the models capability to generalize for uncommon locations.
@@ 6.5. Reinforcement Learning based Location Retrieval
 The MC Dropout Location Retrieval model performs better than the baseline machine learning model when evaluating offline using the mean prediction $\mu$ for each test example with +0.25% in recall and -9.3% in retrieval size. We also qualitatively verified whether the MC Dropout Location Retrieval model explores more for locations with little prior data versus locations that are often searched and booked. For example, the difference between the mean bounds and UCB bounds for San Francisco, California is very small. However, the difference is much larger for Smith Mountain Lake, Virginia that we have little prior data for.
 The MC Dropout Location Retrieval Model was also tested in an online A/B experiment against the baseline production machine learning model retrained on data up to the start of the experiment and the production model trained on data up to 8 months before the start of the experiment. There was no statistically significant difference in performance between the production and retrained production models. The MC Dropout model showed a statistically significant increase of +0.26% in uncancelled bookers compared to the retrained production model, which was primarily driven by surfacing more bookable, affordable, and diverse inventory to guests. This impact is a notable improvement when considering the significant previous iterations and maturity of the system. It’s also comparable to the impact of most published ranking techniques at Airbnb. It now serves all production location search traffic on Airbnb. Traditional multi-armed bandit regret analysis is infeasible for a nontraditional recall maximizing application but model confidence for each prediction offset has increased by 7.33% on average since launch.
@@ 7. Conclusion
We presented how location retrieval was built from the ground up in Airbnb search by solving cold start, generalization, and algorithmic bias. The system evolved from heuristics all the way to the application of reinforcement learning with significant cumulative impact of +2.4% in uncancelled bookers. This is comparable or more impactful than many previously published machine learning techniques within Airbnb search such as (abdool2020managing) (haldar2019applying) (haldar2020improving) (tan2023optimizing) (haldar2023learning). We outlined the successes and failures along the way and the motivation for more advanced techniques to tackle the aforementioned challenges. The Monte Carlo Dropout model, the application of reinforcement learning, proved that it performed better than previous baselines offline and online for business metrics and now serves all production traffic at Airbnb. For future work, there is opportunity to advance the model’s understanding with more complex features. There is also an opportunity to reformulate the retrieval mechanism from retrieval bounds to map cells. This would allow for organically learning booking probabilities for each cell instead of training with hand tuned weights.